OpenStudy (anonymous):

What is the relationship between sin(90°) and sin(−90°)

OpenStudy (freckles):

hint: sin(-x)=-sin(x) since sin(x) is an odd function

OpenStudy (freckles):

Are you there?

OpenStudy (anonymous):

yeah im looking at the question sorry

OpenStudy (freckles):

Here is an example: Take f(x)=x^3 f is an odd function since (-x)^3=-x^3 Consider f(2) and f(-2). What is f(2) and f(-2) equal to?

OpenStudy (anonymous):

so you would plug in the -2 and 2 for x?

OpenStudy (freckles):

yep and what would you get for both

OpenStudy (freckles):

f(2)=2^3=? f(-2)=(-2)^3=?

OpenStudy (anonymous):

\[(2)=2^{3}\] would equal 9?

OpenStudy (freckles):

2*2*2=4*2=8 f(2)=2^3=8

OpenStudy (anonymous):

oohhhh i see lol

OpenStudy (freckles):

f(2)=(2)^3=(2)(2)(2)=(4)(2)=8 f(-2)=(-2)^3=(-2)(-2)(-2)=(4)(-2)=-8 what do you notice about the values 8 and -8?

OpenStudy (freckles):

what do we call that relationship

OpenStudy (freckles):

they are ____ of one another

OpenStudy (anonymous):

similar?

OpenStudy (freckles):

what is the opposite of 8?

OpenStudy (anonymous):

-8

OpenStudy (freckles):

yes they are opposites of one another

OpenStudy (freckles):

since sin(-x)=-sin(x) sin(-x) is the opposite of sin(x) when x doesn't equal n*pi where n is an integer

OpenStudy (anonymous):

so there equal to each other?

OpenStudy (freckles):

sin(-x)=-sin(x) but I didn't say sin(-x) equals sin(x)

OpenStudy (anonymous):

so its greater than?

OpenStudy (freckles):

is that the relationship you want is greater than or less than? because the first thing I thought of is sin(x) is odd so sin(90 deg) and sin(-90 deg) are opposites

OpenStudy (freckles):

you can use your unit circle to determine which outputs 1 and which outputs -1

OpenStudy (anonymous):

sin(90°) < sin(−90°) sin(90°) > sin(−90°) sin(90°) = sin(−90°) cannot be determined

OpenStudy (freckles):

do you know how to use the unit circle?

OpenStudy (anonymous):

kinda

OpenStudy (freckles):

can you find 90 deg on there?

OpenStudy (freckles):

what is the y value coordinate at 90 deg?

OpenStudy (anonymous):

1

OpenStudy (freckles):

right and I already gave you that sin(90) and sin(-90) are opposite values since sin(x) is an odd function

OpenStudy (freckles):

so sin(-90)=?

OpenStudy (anonymous):

1?

OpenStudy (freckles):

the opposite doesn't mean the same as

OpenStudy (anonymous):

-1

OpenStudy (freckles):

yes

OpenStudy (freckles):

sin(-90)=-1 sin(90)=1

OpenStudy (freckles):

can you tell which is greater now?

OpenStudy (anonymous):

would it be the positive?

OpenStudy (freckles):

1 is definitely greater than -1

OpenStudy (anonymous):

sin(90°) < sin(−90°) sin(90°) > sin(−90°) sin(90°) = sin(−90°) cannot be determined

OpenStudy (freckles):

i guess you posted that because you still don't see which is greater sin(90)=1 ___________ -1=sin(-90) I'm trying to get you to see 1>-1 like having 1 dollar is more than having a debt of 1 dollar

OpenStudy (anonymous):

so it would be b

OpenStudy (freckles):

yes because 1>-1

OpenStudy (anonymous):

ok thanks

OpenStudy (freckles):

the first choice says 1<-1 the second choice says 1>-1 the third choice says 1=-1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok thank you